Suppose the random variable Y has a normal distribution with an expected value equal to 12 and variance equal to 16. What is the value of k such as that P(X<k)= 0.25


1
Expert's answer
2022-05-09T15:57:04-0400
"P(X<k)=P(Z<\dfrac{k-12}{\sqrt{16}})=0.25"

"\dfrac{k-12}{4}\approx-0.67449"

"k\approx9.30204"